Plese solve the following integral
∫|x| dx
Limit is from -1 to 2.

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:Hello student, please find answer to your question
I = \int_{-1}^{2}|x|dx
I = \int_{-1}^{0}|x|dx+\int_{0}^{2}|x|dx
I = \int_{-1}^{0}-xdx+\int_{0}^{2}xdx
I = (\frac{-x^{2}}{2})_{-1}^{0} + (\frac{x^{2}}{2})_{0}^{2}
I = \frac{1}{2} + 2
I = \frac{5}{2}